# SCEP Collaboration The projected sensitivity of SCEP experiment to Magnetic Monopole

### v2

@@ III.3 Detection
![Refer to caption](x7.png)
![Refer to caption](x8.png)
![Refer to caption](x9.png)
The readout devices exhibit diverse response characteristics to induction signals, owing to their distinct intrinsic mechanisms. The magnetometer relies on atomic precession and typically demonstrates a response timescale ranging from several tens of microseconds to milliseconds. The complex response function of the magnetometer is commonly modeled in the form of a Lorentzian distribution:
where , with the Bohr magnetic moment and reduced Planck constant, is gyromagnetic ratio of atom caesium and is the spin relaxation time which can be measured experimentally. The mean resonant frequency of the Lorentzian response function of the magnetometer varies depending on the applied static magnetic field along the axis within the gas chamber. In the case of using an ADC for readout, the response function can be simplified and approximated as a constant which is dependent on the gain of the OPA and the input impedance of the ADC, within the bandwidth of interest. In order to reduce spectral waveform distortions due to limited-length time window, specific-shaped time windows, such as the Hamming window [oppenheim1999discrete], are introduced in signal and noise processing.
@@ III.4 Reconstruction and thermal noise
where is the power spectral density of the noise on the induction coil. is the power spectral density of the noise generated during the signal shaping and detection, while represents the product of the conjugates of all response functions present in the same progress. In the ADC-readout scenario, corresponds to the conjugate of the circuit response function , and the is mainly influenced by the noise from the OPA. On the other hand, in the magnetometer-readout scenario, represents the combined conjugate response of both the circuit and magnetometer , and accounts mainly for the thermal noise from the Helmholtz coil. The noise from the magnetometer is negligible. Thermal noise originating from the induction coil significantly influences the overall noise characteristics, especially in the magnetometer-readout scenario. This noise is modeled as Johnson-Nyquist noise [Johnson1928] 222Note the expression of follows the convension in the field of signal processing [somalwar1988transient].:
where is the Boltzmann constant, and is the temperature. It is essential to emphasize that the thermal noise in this particular scenario does not exhibit the characteristic of “white” noise, which is typically assumed to have a frequency-independent power spectral density. Due to the presence of a non-trivial alternating resistance in the induction coil, the thermal noise power increases with higher frequencies. This frequency-dependent behavior of the noise is an important consideration in the analysis and characterization of the system. Fig. 7 shows some waveform examples before and after applying OF for both the ADC-readout and magnetometer-readout scenarios.
The typical SNR is calculated under the assumption of MM velocity being 10-5 light speed, passes perpendicularly through the coil center (denoted as SNR0 in the text). The SNR0s of each prototype coil can be found in Tab. 1. It is worth noticing that the SNR0 does depend on the MM’s speed. SNR0 increases with the increase in MM speed. However, SNR0 gradually tends toward saturation as the effect of alternating resistance increasingly becomes significant. Fig. 8 displays the SNR0 of the prototype induction coils in both ADC-readout and magnetometer-readout modes.
@@ IV Validation of the signal simulation
A validation test is performed to assess the accuracy and reliability of the signal simulation framework. The test mainly aims to validate the waveform amplitudes and shapes of MM signal and noise. These characteristics of the signal waveforms are crucial for predicting the sensitivity of such detection to single Dirac MM.
@@ IV.1 Signal validation
The MM signal validation involves the utilization of a long-thin stimulation coil to generate a pulsed magnetic field that emulates an MM signal on the manufactured induction coil prototypes. Table 1 shows the geometrical parameters of the induction coils, including the wire type, wire diameter, coil minimal/maximal radii, and turn number. Their alternating resistivities are displayed in Fig. 5. The corresponding SNR0s to single Dirac MM are given in Table 1 as well. The highest SNR with ADC readout is about 0.57, mainly limited by the thermal noise of the induction coil and the noise of OPA. On the contrary, the highest SNR0 with magnetometer readout can reach 1.92 because of the low noise level of the magnetometer. However, the parameters of the Helmholtz coils need to be carefully designed. The stimulation coil utilized in the validation has a length of 50 cm and a diameter of 10 mm. The turns number density amounts to approximately 100 per centimeter. During the testing, the stimulation coil passes through the center of the induction coil, perpendicular to its coil plane. The diagram of the induction coil, the stimulation coil, and their positioning are shown in Fig. 9.
![Refer to caption](x10.png)
@@
Due to the prevalence of electromagnetic noise in the surrounding environment and the limited precision of the pulse generator, it is not practically feasible to accurately simulate and test the signal response of the induction coil to a single Dirac MM. In our experimental setup, we generate a voltage pulse with a square wave shape using a pulse generator, and then feed this voltage pulse to the stimulation coil. A resistor with a resistance of 19.36 is connected in series with the stimulation coil. The voltage drop across this resistor is monitored using a digitizer with a sampling rate of 2 MHz, which is connected in parallel to the resistor. This allows us to precisely model the microcurrent passing through the stimulation coil. It should be noted that in our experimental setup, we assume there are no leak fields associated with the tightly wound stimulation coil. By employing this stimulation process, we are able to generate magnetic flux pulses on the induction coil that closely mimic those produced by the passing of MM in temporal shape.
@@
Such test is performed for all three prototype induction coils with the ADC-readout scenario. For the magnetometer-readout scenario, V2 coil is tested which is expected to have the largest SNR among all three prototype coils. Fig. 10 shows the comparison between the measured and predicted test signals in the time domain. The readout signals can be parameterized as:
@@
where is the phase. The , , and are the amplitude, frequency, and decay rate, respectively. These three parameters are compared between the expectation and measurement. The results of the comparison are summarized in Table 2. The measured frequencies and decay rates are consistent with the predictions, with bias no more than 0.3% and 8.7% for the frequency and decay rate, respectively. This validates our response function models of the circuit and magnetometer. The largest amplitude differences observed between measurements and predictions are about 12.5% for ADC readout and 9.2% for magnetometer readout. This is considered to be due to the leak fields and non-even turn density of the stimulation coil. Particularly the field leakage is more severe since the size of V3 coil is the largest among the tested ones. In addition, the lower amplitude seen in the measurement with magnetometer readout could be due to the potential bias of the effective Lorentzian response shown in Eq. 6.
@@ IV.2 Noise validation
To determine the intrinsic thermal noise power spectrum, the V2 coil is enclosed within a grounded metal box constructed of copper, which served as a Faraday cage. Fig. 11 displays the power spectra of the V2 coil under two conditions: when the coil is exposed to air and when it is sealed inside the copper box. A significant reduction in noise is observed when the coil is enclosed in the copper box, indicating the presence of a strong electromagnetic noise background in the air. Furthermore, the frequency domain analysis revealed distinct peak-like structures upon placing the coil inside the copper box. These peaks corresponded to multiples of the common frequency in utility, suggesting the presence of leaked-in electromagnetic waves within the copper box, likely originating from the signal connectors. This hypothesis is supported by the observation that the orientation of the induction coil influences the level of noise detected. The lowest noise level is observed when the coil axis is in a vertical position, as shown in Fig. 11. The observed noise frequency spectrum closely resembled the predicted one by the simulation, with a slightly lower amplitude (8.0%) at the resonant frequency.
![Refer to caption](x12.png)
V Projected Sensitivity to Magnetic Monopole
@@
The search for MMs eventually will be conducted using an array of induction coils. The top and bottom of the coil array are equipped with plastic scintillators, as depicted in Fig. 1. The experimental setup can be situated either on Earth or in deep space, such as on the Moon. In both scenarios, the detector array will be exposed to significant levels of background particles, specifically protons, muons, and alpha particles (helium nuclei). These background particles leave scintillation signals in the scintillators and can accidentally pile-up with the thermal noise in the induction coils, creating false MM signals. The impact of this background can be mitigated by requiring more layers of the induction coils and the particle detectors.
@@
To assess the sensitivity of the detector array to MMs, we employ a simple ideal benchmark configuration. This configuration consists of induction coils with a diameter of 12 cm (same as V2 coil), arranged vertically and compactly instrumented. The array’s size is assumed to be sufficiently large to disregard any edge effects. The alternating resistance (equivalent to the thermal noise configuration) of each induction coil is assumed to followed the one of V2 coil, and each coil is assumed to have negligible height. In this benchmark analysis, a simple over-threshold trigger is conducted on waveform of each induction coil after the OF applied. The coil array is equipped with PS layers at the top and bottom, and these layers are positioned approximately 1 meter apart in the benchmark. Each PS layer is composed of two sets of PS panels arranged perpendicular to each other. This benchmark arrangement allows for the reconstruction of events’ transverse positions.
V.1 Acceptance to GUT-MM induction
@@
The GUT-MM is assumed to exhibit isotropic behavior in terrestrial and deep-space environments. However, due to the round geometry of the induction coil, there is an inherent acceptance loss of (1-/4) for each layer of coils. We consider simply the coil layers are identical and sufficiently close to each other, so that we can consider such benchmark setup having a conservative acceptance loss of (1-/4) due to coil geometry. Optimizing the coil geometry and arrangement between layers can alleviate the acceptance loss to some extent.
@@
The dependence of the SNR on the point of MM transpassing the coil is weak. In Fig. 12, the average acceptance to GUT-MM is displayed as a function of the transverse angle (), considering various assumptions regarding the SNR0. Only when the approaches , the acceptance drops quickly. The Fig. 13 shows the angle-averaged acceptance as a function of . All calculations are based on an MM speed of 10-5 times the speed of light.
![Refer to caption](x13.png)
![Refer to caption](x14.png)
@@ V.2 Background of Induction Signal
The cosmic rays and their secondaries, such as high-energy protons, muons, and electrons, that are common in terrestrial and deep-space environments, deposit energy in the top and bottom PSs, but produce negligible induction signals. These particles possess magnetic dipoles and travel mostly at relativistic speeds, resulting in a distinct induction pulse shape with a resonant shape and faster time response compared to those from the MMs. Therefore, we consider that relativistic cosmic rays and their secondaries do not produce any significant direct background in the induction signals.
@@
However, the energy depositions detected by the PSs may coincide with the abundant thermal noise present in the induction coils, leading to mis-identified MM signals. As discussed in Subsection III.4, the thermal noise arises from the non-zero alternating resistances of the induction coil and constitutes the main background for the MM induction signal search. In this analysis, a simple over-threshold approach is employed as the trigger method on a single coil. The dependence of the acceptance and noise rate on the threshold (denoted as in the text) with different assumed are shown in the Fig. 13. The rate of the mis-triggering of thermal noise and the acceptance across the coils to form a track-like event can be expressed as follows:
@@
Here, and represent the mis-triggered noise rate and acceptance of a single induction coil at a given threshold, denotes the number of coils required to detect the induction trigger (coincidence number), and represents the time response of the induction signal, which is related to the resonant frequency of the induction coil. For the benchmark analysis, we assume .
V.3 Background with Particle Coincidence
![Refer to caption](x15.png)
![Refer to caption](x16.png)
![Refer to caption](x17.png)
![Refer to caption](x18.png)
![Refer to caption](x19.png)
@@ V.3 Background with Particle Coincidence
The rate of reconstructed scintillation signals on PSs is mainly affected by two factors: random pileups occurring between the top and bottom PSs, and the passage of a relativistic particle through both PSs. The energy threshold of PS is estimated to be 0.1 MeV, below which the contribution of SiPM dark count pileup rocket-rises. This reconstructed scintillation signal necessitates the presence of two energy depositions, one on each of the top and bottom PSs. It is crucial for the reconstructed energies, timings, and transverse positions of these two energy depositions to align with the expected energy, time of flight (ToF), and track characteristics of the MM of interest. The differential reconstructed scintillation signal rate per unit area per radian on the two PS panels can be expressed as the sum of two components: the rate arising from pileup events, denoted as , and the rate resulting from direct passage of particles, denoted as . Both contributions are related to the effective scintillation rate on a single PS given a zenith angle, denoted as in unit of cm-2s-1sr-1:
@@
where is the pileup time window determined by the PS time resolution, which is assumed to be 10 ns [zhou2020time], and =1 m is the distance between top and bottom PSs. We require to be the rate after an energy range cut that covers 99.5% (3) of MMs. Such cuts give an effective efficiency to background particles of , which also applies to direct particle passage but with two layers of PSs both requiring such energy selection. The is the particle flux. Note that depends on the speed of MM. The total background rate that takes all the angles into account can be expressed as:
@@
In practice, the value of is influenced by several factors, including the background particle flux and spectrum, the ability to determine the direction of the MM using induction signals on the coils, and the energy resolution of the PS. Also, in order to determine the , the amount of light produced in the PS by Dirac MM is needed, which depends on the MM velocity. The detectable stopping power, also known as the light yield, of Dirac MM on the PS as a function of MM speed is presented in the top panel of Fig. 14, based on the calculations in Ref. [derkaoui1999energy]. To differentiate the energy deposition of the MM from common background particles such as protons, electrons, alpha particles, and muons, we require that the reconstructed energy falls within 3 times the energy resolution. The intrinsic energy resolution of the PS, as a function of the total deposited energy, is obtained through optical simulation using GEANT4 [agostinelli2003geant4]. The energy resolution is illustrated in Fig.15. The reconstruction resolution of the transverse position in the PS-based array primarily depends on the width of the PS panel, which is significantly smaller than the size of the induction coil. Consequently, the track reconstructed by the PS exhibits much higher resolution compared to the one reconstructed by the induction coils. For this benchmark analysis, we conservatively consider after the coincidence requirement to be the background rate within a 12 cm-diameter circle, which corresponds to the size of the V2 coil used in the estimation.
@@
In a terrestrial detector situated at sea level, the primary background particles are atmospheric muons, as well as the residual high-energy protons. Muons with kinetic energies ranging from hundreds of MeV to hundreds of GeV exhibit minimal ionizing behavior when interacting with matter, enabling them to easily traverse the surrounding materials near the detector, including the top and bottom PSs. On the other hand, the proton flux experiences a significant reduction as it traverses the atmosphere due to ionization and radiative processes. However, protons leave a higher ionization density in the PS compared to muons, approaching the ionization density that could be produced by Dirac MMs within a specific range of speeds. The stopping powers of protons and muons, corresponding to detectable energy ranges, in the PS are calculated based on the PSTAR and ASTAR databases [international1993stopping], PDG sources [groom2001muon], and the methodology outlined in Ref. [derkaoui1999energy]. These stopping powers are presented in Fig.14. To model the flux and angular distribution of atmospheric muons at sea level, the Bugaev/Reyna model [su2021comparison] is employed, while the flux of high-energy protons is simulated using CRY algorithms [hagmann2007cosmic]. The fluxes can be observed in Fig.16.
On Moon, the muons are no longer dominant because of the absence of atmosphere. In deep space, high-energy protons and helium nuclei emerge as the prevailing particles, as evidenced by findings from the AMS [aguilar2021periodicities, aguilar2022properties], DAMPE [dampe2019measurement, dampe2022detection], and CALET [marrocchesi2019measurement, brogi2022measurement] experiments. Assuming the negligible influence of Earth’s magnetic field on the Moon, it is conservatively assumed that the fluxes of protons, helium nuclei, carbon nuclei, and iron nuclei are homogeneous in all direction. The fluxes as a function of particle kinetic energies are taken from Ref. [waller2020simulated], and are also shown in Fig. 16.
@@
A toy Monte Carlo (MC) simulation is performed to calculate the background and MM spectra. The spread in deposited energy caused by varying travel lengths inside the PS due to different incoming particle angles is also taken into consideration in the toy MC. In the final analysis of the top and bottom PSs, we are able to provide the reconstructed zenith angles of incoming background particles or MMs or the “fake” particles reconstructed from pileups. Therefore, all the deposit energies are corrected to the equivalent deposit energy as if it passes through the material perpendicularly, denoted as . The top panels of Fig.17 illustrate the predicted deposit energy spectra of Dirac MMs and background particles in a single PS layer (before correction). The middle panels give the differential rates for a single PS layer. The lower panels depict the total scintillation background rate for different assumed MM speeds, requiring that we select an range that covers 99.5% (3) of MMs and also covers 99.5% (3) of MMs’ time of flight. Considering the energy threshold of 0.1 MeV in the plastic scintillator, we do not expect its acceptance of MMs with speeds lower than about 2.510-4 light speed.

### v3

@@ III.3 Detection
The readout devices exhibit diverse response characteristics to the induction signals, owing to their distinct intrinsic mechanisms. The magnetometer relies on the atomic precession and typically demonstrates a response timescale ranging from several tens of microseconds to several milliseconds. The complex response function of the magnetometer is commonly modeled in the form of a Lorentzian distribution:
where , with being the Bohr magnetic moment and the reduced Planck constant, is the gyromagnetic ratio of the target atom and is the spin relaxation time which can be measured experimentally. The mean resonant frequency of the Lorentzian response function of the magnetometer depends on the applied static magnetic field along the axis within the gas chamber. In the case of the ADC direct readout, the response function can be simplified and approximated as a constant which is dependent on the gain of the OPA and the input impedance of the ADC, within the bandwidth of interest.
@@ III.4 Reconstruction and thermal noise
where is the power spectral density of the noise on the induction coil. is the power spectral density of the noise generated during the signal shaping and detection, while represents the product of the conjugates of all response functions present in the readout process. In the ADC-readout scenario, corresponds to the conjugate of the circuit response function , and the is mainly influenced by the noise from the OPA. In the magnetometer-readout scenario, represents the combined conjugate response of both the circuit and magnetometer , and accounts mainly for the thermal noise from the Helmholtz coil. The noise from the magnetometer is negligible. Thermal noise originating from the induction coil significantly influences the overall noise characteristics, especially in the magnetometer-readout scenario. This noise is modeled as Johnson-Nyquist noise [Johnson1928] 222Note the expression of follows the convension in the field of signal processing [somalwar1988transient].:
where is the Boltzmann constant, and is the temperature. It is essential to emphasize that the thermal noise in this particular scenario does not exhibit the characteristic of “white” noise, which is typically assumed to have a frequency-independent power spectral density. Due to the presence of a non-trivial alternating resistance in the induction coil, the thermal noise power increases with higher frequencies. Fig. 7 shows some waveform examples before and after applying OF for both the ADC-readout and magnetometer-readout scenarios. In order to reduce the spectral waveform distortions due to the limited-length time window, specific-shaped time windows, such as the Hamming window [oppenheim1999discrete], are introduced in the signal and noise processing.
The typical SNR is calculated under the assumption that the MM passes perpendicularly through the coil center with the speed of 10-5 light speed (denoted as SNR0 in the text). After optimizing the circuit configuration for each test coil, the SNR0s of each coil can be found in Table 1. It is worth noting that the SNR depends on the MM’s speed. Fig. 8 displays the SNR’s velocity dependence of the prototype induction coils in both ADC-readout and magnetometer-readout modes. The SNR increases with the increase in MM speed. However, SNR gradually tends toward saturation because the alternating resistance increasingly becomes significant at high frequency and the resonant frequency is preliminarily set to about several kHz in this study. Increasing the resonant frequencies can potentially improve the SNR at the high velocity ranges.
@@ IV Validation of the signal simulation
A validation test is performed to assess the accuracy and reliability of the signal simulation framework. The test mainly aims to validate the waveform amplitudes and shapes of the MM signal and the thermal noise. These characteristics of the signal waveforms basically define the SNR0 and are crucial in the prediction of the sensitivity to the single Dirac MM.
@@ IV.1 Signal validation
The MM signal validation involves the utilization of a long-thin stimulation coil to generate a pulsed magnetic field that emulates an induction signal that could be caused by large numbers of Dirac MMs on the test coils. The highest SNR0 among the three test coils with ADC readout is about 0.57, mainly limited by the thermal noise of the induction coil and the noise of OPA, as shown in Table 1. On the contrary, the highest SNR0 with magnetometer readout can reach 1.92 because of the low noise level of the magnetometer. The stimulation coil utilized in the validation has a length of 50 cm and a diameter of 10 mm. The turns number density amounts to approximately 100 per centimeter. During the testing, the stimulation coil passes through the center of the induction coil, perpendicular to its coil plane.
@@
In our experimental setup, we generate a voltage pulse with a square wave shape using a pulse generator. Due to the prevalence of the electromagnetic noise in the surrounding environment and the limited precision of the pulse generator, it is not practically feasible to accurately emulate and test the signal response to a single Dirac MM. The generated magnetic flux is approximately 1.710-8 Wb, equivalent to the flux created by about 4M Dirac MMs. A resistor with a resistance of about 20 is connected in series with the stimulation coil. The voltage drop across this resistor is monitored using a digitizer with a sampling rate of 2 MHz, which is connected in parallel to the resistor. This allows us to model the microcurrent passing through the stimulation coil. It should be noted that in our experimental setup, we assume there are no leak fields associated with the tightly wound stimulation coil.
@@
Such test is performed for all three test induction coils with the ADC-readout scenario. For the magnetometer-readout scenario, the V2 coil is tested which is expected to have the largest SNR among all three coils. Fig. 9 shows the comparison between the measured and predicted test signals in the time domain. The readout signals can be parameterized as:
@@
where the , , , and are the amplitude, frequency, decay rate, and phase, respectively. The , , and are compared between the expectation and the measurement. The results of the comparison are summarized in Table 2. The measured frequencies and decay rates are consistent with the predictions, with bias no more than 0.3% and 8.7%, respectively. This validates our response function models of the circuit and magnetometer. The largest amplitude differences observed between measurements and predictions are about 12.5% for the ADC readout and 9.2% for the magnetometer readout. This is considered to be likely due to the leak fields and uneven turn density of the stimulation coil. Particularly the field leakage is more severe for V3 coil since the size of V3 coil is the largest among the tested ones. In addition, the lower amplitude seen in the measurement with the magnetometer readout could be also due to the potential bias of the effective Lorentzian response shown in Eq. 6.
@@ IV.2 Noise validation
To determine the intrinsic thermal noise power spectrum, V1 coil is enclosed within a grounded metal box constructed of copper, which served as a Faraday cage. Fig. 10 displays the power spectra of V1 coil under two conditions: when the coil is exposed to the air and when it is sealed inside the grounded copper box. A significant reduction in noise is observed when the coil is enclosed in the copper box, indicating the presence of a strong electromagnetic noise background in the air. Furthermore, the frequency domain analysis revealed distinct peak-like structures upon placing the coil inside the copper box. These peaks corresponded to the multiples of the common frequency in utility, suggesting the presence of the leaked-in electromagnetic waves within the copper box, likely originating from the signal connectors. This hypothesis is supported by the observation that the orientation of the induction coil influences the level of noise detected. The lowest noise level is observed when the coil axis is in a vertical position, as shown in Fig. 10. The observed noise frequency spectrum closely resembled the predicted one by the simulation, with a slightly lower amplitude (8.0%) at the resonant frequency.
V Sensitivity to Cosmic Magnetic Monopole
@@
The search for the MMs will be conducted using an array of induction coils. The top and bottom of the coil array will be equipped with PSs, as depicted in Fig. 1. We take the detector situated at the sea level on Earth as the benchmark. The dominant background is the pileup between the scintillation signals caused by the cosmic ray induced background particles (mainly muons and protons) in the PSs and the thermal noise in the induction coils. The impact of this background can be mitigated by requiring more layers of the induction coils and the particle detectors.
@@
To assess the sensitivity of the detector array to the cosmic MMs, we employ a simple ideal configuration. This configuration consists of induction coils with a diameter of 12 cm (same as V2 coil), arranged vertically and compactly instrumented. The array’s size is assumed to be sufficiently large to disregard any edge effects. The alternating resistance’s frequency dependence of each induction coil is assumed to followed the one of V2 coil, and each coil is assumed to have negligible height. In this analysis, a simple over-threshold trigger is conducted on waveform of each induction coil after the OF applied. The coil array is equipped with the PS layers at the top and bottom, and these layers are positioned approximately 1 meter apart. Each PS layer is composed of two sets of PS panels arranged perpendicular to each other. This arrangement allows for the reconstruction of events’ transverse positions.
V.1 Acceptance to cosmic MM induction
@@
The cosmic MM is assumed to exhibit isotropic behavior. However, due to the geometry of the induction coil, there is an inherent acceptance loss of (1-/4) for each layer of coils. We consider simply the coil layers are identical and sufficiently close to each other, so that we can consider such setup having a conservative acceptance loss of (1-/4) due to coil geometry. Optimizing the coil geometry and arrangement between layers can alleviate the acceptance loss to some extent.
@@
The dependence of the SNR on the transverse distance of the MM to the coil center is weak. In Fig. 11, the average acceptance to cosmic MM is displayed as a function of the polar angle (), considering various assumptions regarding the SNR0. Only when the approaches , the acceptance drops quickly.
@@ V.2 Background of Induction Signal
The cosmic rays and their secondaries, such as high-energy protons, muons, and electrons, that are common in the terrestrial environment, deposit energy in the top and bottom PSs, but produce negligible induction signals. These particles possess magnetic dipoles and travel mostly at relativistic speeds, resulting in a distinct resonant induction pulse shape and faster time response compared to those from the MMs. Therefore, we consider that relativistic cosmic rays and their secondaries do not produce any significant background for the induction signals.
@@
However, the energy depositions detected by the PSs may coincide with the abundant thermal noise present in the induction coils. As discussed in Subsection III.4, the thermal noise arises from the non-zero alternating resistances of the induction coil and constitutes the main background for the MM induction signal search. In this analysis, a simple over-threshold approach is employed as the trigger method on a single coil. The dependence of the acceptance and noise rate on the threshold (denoted as in the text) with different assumed s are shown in Fig. 12. The rate of the mis-triggering of thermal noise and the acceptance across the coils to form a track-like event can be expressed as follows:
@@
Here, and represent the mis-triggered noise rate and acceptance of a single induction coil at a given threshold . denotes the number of coils required to detect the induction trigger (coincidence number). represents the time response of the induction signal, which is related to the resonant frequency of the induction coil. For the analysis, we assume .
V.3 Background of Scintillation Signal
@@
The rate of the reconstructed scintillation signals on the PSs is mainly affected by two factors: random pileups occurring between the top and bottom PSs, and the passage of a relativistic particle through both PSs. The energy threshold of PS is estimated to be 0.1 MeV, below which the contribution of SiPM dark count pileup skyrockets. This reconstructed scintillation signal necessitates the presence of two energy depositions, one on each of the top and bottom PSs. It is crucial for the reconstructed energies, timings, and transverse positions of these two energy depositions to align with the expected energy, time of flight (ToF), and track characteristics of the MM of interest. The differential reconstructed scintillation signal rate per unit area per radian on the two PS panels can be expressed as the sum of two components: the rate arising from pileup events, denoted as , and the rate resulting from direct passage of particles, denoted as .
@@
where is the pileup time window determined by the PS time resolution, which is assumed to be 10 ns [zhou2020time], and =1 m is the distance between top and bottom PSs. is the effective scintillation rate on a single PS layer given a zenith angle. , , and are in unit of cm-2s-1sr-1. We require to be the rate after an energy range cut that has a coverage of 99.5% (3) for the MMs. Such cuts give an effective efficiency to background particles of , which also applies to direct particle passage but with two layers of PSs both requiring such energy selection. The is the particle flux. Note that depends on the speed of MM. The total scintillation background rate per unit area that takes all the angles into account can be expressed as:
@@
In practice, the value of is influenced by several factors, including the background particle flux and spectrum, the ability to determine the direction of the MM using the induction signals, and the energy resolution of the PS. Also, in order to determine the , the amount of light produced in the PS by a Dirac MM is needed, which depends on the MM velocity. The detectable stopping power, also known as the light yield, of a Dirac MM on the PS as a function of MM speed is presented in the top panel of Fig. 13, based on the calculations in Ref. [derkaoui1999energy]. To differentiate the energy deposition of the MM from common background particles such as protons, electrons, alpha particles, and muons, we require that the reconstructed energy falls within 3 times the energy resolution. The intrinsic energy resolution of the PS, as a function of the total deposited energy, is obtained through an optical simulation based on GEANT4 [agostinelli2003geant4]. The energy resolution is illustrated in Fig.14. The reconstruction resolution of the transverse position in the PS-based array primarily depends on the width of the PS panel, which is significantly smaller than the size of the induction coil. Consequently, the track reconstructed by the PS exhibits much higher resolution compared to the one reconstructed by the induction coils. For this analysis, we conservatively consider after the coincidence requirement to be the background rate within a 12 cm-diameter circle, which corresponds to the size of the V2 coil used in the estimation.
@@
In a terrestrial detector situated at sea level, the primary background particles are the atmospheric muons, as well as the residual high-energy protons. Muons with kinetic energies ranging from hundreds of MeV to hundreds of GeV exhibit minimal ionizing behavior when interacting with matter, enabling them to easily traverse the surrounding materials near the detector, including the top and bottom PSs. On the other hand, the proton flux experiences a significant reduction as it traverses the atmosphere due to the ionization and radiative processes. The protons leave a higher ionization density in the PS compared to the muons, approaching the ionization density that could be produced by the Dirac MMs within a certain range of speeds. The stopping power and the light yields of protons and muons, corresponding to detectable energy ranges, in the PS are calculated based on the PSTAR and ASTAR databases [international1993stopping], PDG sources [groom2001muon], and the methodology outlined in Ref. [derkaoui1999energy]. These light yields are presented in Fig.13. To model the flux and angular distribution of atmospheric muons at sea level, the Bugaev/Reyna model [su2021comparison] is employed, while the flux of high-energy protons is simulated using CRY algorithms [hagmann2007cosmic]. The fluxes are shown in Fig.15.
@@
We performed a toy Monte Carlo (MC) simulation to calculate the background and MM energy deposition spectra. The simulation takes into account the spread in the deposited energy caused by the variation in particle traveling lengths within the PS due to different incoming particle angles. In the final analysis of the top and bottom PSs, we reconstruct the zenith angles of the incoming background particles, MMs, or ”fake” particles reconstructed from pileup events. All the deposited energies are corrected to the equivalent energy deposition when the particle passed through the material perpendicularly, denoted as . The top panel of Fig. 16 illustrates the predicted deposited energy spectra of Dirac MMs with different assumed speeds and the background particles in a single PS layer (before correction). The middle panel gives the differential rates for a single PS layer. The lower panel depicts the total scintillation background rate for different assumed MM speeds, requiring that we select an range that covers 99.5% (3) of the MMs and a ToF that also covers 99.5% (3) of the MMs. Considering the energy threshold of 0.1 MeV in the PS, the acceptance to MMs with speeds lower than about 2.510-4 light speed is negligible.